A 1.5 mm -diameter, 20 cm -long copper wire carries a 4.5 A current.

What is the potential difference between the ends of the wire?

in mV
Answer #1

here,

diameter , d = 1.5 mm

d = 0.75 * 10^-3 m

length of wire , l = 0.2 m

the resistance of wire , r = p * l area

r = 1.68 * 10^-8 * 0.2 /( pi * (0.75 * 10^-3)^2)

r = 0.19 * 10^-2 ohm

current , I = 4.5 A

the potential difference between the ends of the wire , V = I * R

V = 4.5 * 0.19 * 10^-2 V

V = 8.56 * 10^-3 V

V = 8.56 mV